# Maria Bazotte, Margarida Carvalho, and Thibaut Vidal

### v3

@@
[…]
Solving Two-Stage Stochastic Programs with Endogenous Uncertainty via Random Variable Transformation
[…]
Maria Bazotte
@@
[…]
Real-world decision-making problems often involve decision-dependent uncertainty, where the probability distribution of the random vector depends on the model decisions. Few studies focus on two-stage stochastic programs with this type of endogenous uncertainty, and those that do lack general methodologies. We propose a general method for solving a class of these programs based on random variable transformation, a technique widely employed in probability and statistics. The random variable transformation converts a stochastic program with endogenous uncertainty (original program) into an equivalent stochastic program with decision-independent uncertainty (transformed program), for which solution procedures are well-studied. Additionally, endogenous uncertainty usually leads to nonlinear nonconvex programs, which are theoretically intractable. Nonetheless, we show that, for some classical endogenous distributions, the proposed method yields mixed-integer linear or convex programs with exogenous uncertainty. We validate this method by applying it to a network design and facility-protection problem, considering distinct decision-dependent distributions for the random variables. While the original formulation of this problem is nonlinear nonconvex for most endogenous distributions, the proposed method transforms it into mixed-integer linear programs with exogenous uncertainty. We solve these transformed programs with the sample average approximation method. We highlight the superior performance of our approach compared to solving the original program in the case a mixed-integer linear formulation of this program exists.
@@ 1 Introduction
Uncertainty is an inherent characteristic of most real-world decision-making problems in fields such as network design planning (Contreras et al. 2011), vehicle routing (Gendreau et al. 1996), and scheduling (Restrepo et al. 2017). Stochastic programming is a well-known framework in mathematical programming to model optimization problems with uncertain parameters; see, e.g., (Birge and Louveaux 2011) and (Shapiro et al. 2021) for a summary of the theoretical aspects and the most popular algorithms. Research in this area has focused mainly on exogenous or decision-independent uncertainty (Li and Grossmann 2021), where the random vector is independent of the model decisions. In contrast, endogenous or decision-dependent uncertainty, where the random vector depends on the decisions, has received less attention, although it is prevalent in critical applications such as pre-disaster supply chain management (Peeta et al. 2010, Bhuiyan et al. 2020), production planning (Ekin 2018), and R&D portfolio selection (Solak et al. 2010).
In the context of endogenous uncertainty, different types of decision dependency may occur (Jonsbrten et al. 1998). Goel and Grossmann (2006) proposed classifying decision dependency into two types: For problems with Type 1 endogenous uncertainty or decision-dependent probability, decisions impact the underlying probability distribution of the random parameters. Although little research focuses on this type of uncertainty, as noted by Li and Grossmann (2021), it deals with various problems of practical relevance, such as the design of communication networks, power plants, and warehouses and resource allocation in queuing systems (Pflug 1990). With Type 2 endogenous uncertainty, also called “exogenous uncertainty with endogenous observability” (Mercier and Van Hentenryck 2011), decisions influence the time at which the realization of the random variables is totally or partially revealed. Thus, with Type 2 endogenous uncertainty, the scenario tree depends on the decisions (Jonsbrten et al. 1998). Stochastic programs with Type 2 uncertainty have received more attention than those with Type 1, being mostly multistage programs.  Apap and Grossmann (2017) and Li and Grossmann (2021) provide concise reviews of methodologies and algorithms for multistage programs with Type 2 endogenous uncertainty. In the present work, we focus on Type 1 endogenous uncertainty, in particular, on two-stage stochastic mixed-integer linear programs with decision-dependent probabilities (2-SMIPD), in which the probability distribution depends on the first-stage decisions. In what follows, we refer to Type 1 endogenous uncertainty as endogenous uncertainty.
@@
The main contribution of this work is a general method based on random variable transformation to address 2-SMIPD with discrete or continuous endogenous random variables. This method resolves two main issues in the literature: the lack of generalist methods and the difficulty of solving the usual nonconvex nonlinear programs with endogenous uncertainty. More specifically, the contributions of this work are as follows: (i) We define a transformation function to model the endogenous uncertainty by combining first-stage decisions with decision-independent random variables, thereby converting stochastic programs with endogenous uncertainty (i.e., original programs) into stochastic programs with exogenous uncertainty (i.e., transformed programs). (ii) We identify different techniques to define appropriate transformation functions to deal with endogenous random variables following classical decision-dependent probability distributions, including a commonly studied case known as the “discrete selection of distributions”. (iii) We show that, for these distributions, the obtained stochastic programs with exogenous uncertainty have a convex continuous relaxation. (iv) We investigate the implications of endogenous uncertainty when applying sampling methods, specifically the sample average approximation (SAA) method, to the transformed program. (v) We apply the proposed approach to the network design and facility protection problem (NDFPP), considering different decision-dependent distributions. (vi) Finally, for all endogenous distributions, we analyze the statistical estimators of the SAA applied to the transformed program of the NDFPP. In less than 1 hour on average, the SAA applied to the transformed program finds feasible solutions with significantly small over-estimated optimality gaps (between $-1\%$ and $1\%$) and a high confidence interval for all decision-dependent distributions. For the NDFPP with most endogenous distributions, the deterministic equivalent program (DEP) of the original program is nonlinear and nonconvex. The only exception is the discrete selection of distributions case, for which the DEP of the original program is a mixed-integer linear program (MILP) obtained by auxiliary linearizations (Bhuiyan et al. 2020). In a considerably shorter time, the SAA applied to the transformed program obtains a better over-estimator of the optimality gap than the optimality gap obtained by solving this mixed-integer DEP (a difference of around $24\%$ on average). Solving the transformed program with the SAA also obtains better solutions than when using an average scenario, i.e., the expected value problem (EV).
@@
The remainder of this work is structured as follows: Section 2 formalizes the class 2-SMIPD and discusses related works on endogenous uncertainty. Section 3 presents our general methodology to obtain the transformed program and discusses the application of methods such as the SAA to solve this resulting program. Section 4 catalogs techniques to define transformations to various endogenous distributions, and Section 5 explains how to apply this approach to the NDFPP. Finally, Section 6 reports the experimental results, comparing the proposed method with other solution procedures, and Section 7 concludes our analyses.
@@ 2 Problem Definition and Related Works
We focus on two-stage stochastic mixed-integer linear programs with decision-dependent probabilities (2-SMIPD). Let $x\in\mathbb{R}^{n_{1}}_{+}$ be the vector of first-stage variables (i.e., the “here-and-now” decisions) and $\xi(\omega)\in\mathbb{R}^{U}$ a random vector that depends on the random events $\omega\in\Omega$. For the sake of simplicity, we also denote $\boldsymbol{\xi}=\xi(\omega)$ and designate $\xi$ as a realization of this random vector. We use $\mathbb{E}_{\boldsymbol{\xi}\mid x}[\cdot]$ to denote the expected value operator with respect to $\boldsymbol{\xi}$ given the first-stage variables $x$. Finally, $y(\omega)\in\mathbb{R}^{n_{2}}_{+}$ is the vector of second-stage variables given a realization of the random events. The standard mathematical formulation for the 2-SMIPD is
$\min$ $c^{\top}x+\mathbb{E}_{\boldsymbol{\xi}\mid x}\left[Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}\right]$ (1a)
s.t. $Ax\geq b$ (1b)
$x\in\mathcal{X}$ (1c)
where $\textrm{where }Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}=\min\{q(\omega)^{\top}y(\omega):T(\omega)x+W(\omega)y(\omega)\geq h(\omega),y(\omega)\in\mathcal{Y}\}.$ (1d)
@@
In Program (1), $\mathcal{X}\subseteq\mathbb{R}^{n_{1}}_{+}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n_{2}}_{+}$ are two compact sets imposing non-negativity and continuous, integer, or binary restrictions. The first stage has the matrices $A$, $b$, and $c$ with appropriate dimensions. The second-stage Program (1d) is established after the realization of the random events $\omega\in\Omega$ and has the matrices $q(\omega)$, $h(\omega)$, $T(\omega)$, and $W(\omega)$ with possible random elements and appropriate dimensions. Thus, the random vector $\xi^{T}(\omega)$ combines these random elements, i.e., $\xi^{T}(\omega)=\boldsymbol{(}q(\omega)^{\top},h(\omega)^{\top},T(\omega),W(\omega)\boldsymbol{)}$, and its support (according to its unconditional distribution on $x$) is denoted by $\Xi\subseteq\mathbb{R}^{U}$. For the sake of clarity, we define the random vectors $\boldsymbol{\xi}_{x}^{\top}=\boldsymbol{(}q_{x}(\omega)^{\top},h_{x}(\omega)^{\top},T_{x}(\omega),W_{x}(\omega)\boldsymbol{)}$ for $x\in\mathcal{X}$, which follow the conditional probability distribution of the random vector $\boldsymbol{\xi}$ given $x$ and have support $\Xi_{x}\subseteq\Xi$. Finally, we set $g(x)=\mathbb{E}_{\boldsymbol{\xi}\mid x}[Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}]$.
@@
The 2-SMIPD assumes that the probability distribution of $\boldsymbol{\xi}$ depends on the first-stage decisions $x\in\mathcal{X}$ (i.e., it computes the conditional expectation $\mathbb{E}_{\boldsymbol{\xi}|x}[\cdot]$ instead of the standard expectation $\mathbb{E}_{\boldsymbol{\xi}}[\cdot]$ used in the exogenous case). This endogenous probability influences the properties of the resulting formulation. For example, unlike the exogenous case, the convexity of the continuous relaxation of $Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}$ in Program (1) does not necessarily yield the convexity of the continuous relaxation of $\mathbb{E}_{\boldsymbol{\xi}|x}[Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}]$ (Varaiya and Wets 1989, Dupacová 2006). As a result, algorithms commonly used for solving two-stage programs with exogenous uncertainty may not be suitable for 2-SMIPD.
@@
For the sake of simplicity, we next discuss our reasoning, assuming that the random vector $\boldsymbol{\xi}$ is discrete and has finite support $\Xi$; the process is analogous for the continuous case. We write the DEP in the extensive form of Program (1):
$\min$ $c^{\top}x+\sum_{k\in K}p_{k}(x)q_{k}^{\top}y_{k}$ (2a)
@@
$T_{k}x+W_{k}y_{k}\geq h_{k}$ $\forall\ k\in K$ (2b)
$y_{k}\in\mathcal{Y}$ $\forall\ k\in K,$ (2c)
where $K$ is the set of realizations of $\boldsymbol{\xi}$ (i.e., the set of scenarios defined by the support $\Xi$). For each scenario $k\in K$, $q_{k}$, $h_{k}$, $T_{k}$, and $W_{k}$ have appropriate dimensions, and $\xi_{k}^{\top}=(q_{k}^{\top},h_{k}^{\top},T_{k},W_{k})$. In addition, $p_{k}(x):\mathcal{X}\rightarrow[0,1]$ is a function returning the conditional probability of scenario $k\in K$ given the first-stage variables $x$, in other words, $\boldsymbol{\xi}_{x}$ follows the probability distribution defined by $p_{k}(x)$ for every $k\in K$. We remark that scenarios in $\Xi\backslash\Xi_{x}$ have zero probability according to $p_{k}(x)$.
@@ 2.1 Background of Decision-Dependent Probability Distribution
In Program (2), the probability function $p_{k}(x)$ multiplies the second-stage objective term $q_{k}^{\top}y_{k}$, leading to nonlinear terms. Handling this nonlinearity requires describing $p_{k}(x)$, which depends on the assumptions of $\boldsymbol{\xi}$. When the $U$ elements of the random vector $\boldsymbol{\xi}$ are independent, we have
$p_{k}(x)=\prod_{u=1}^{U}\hat{p}_{\hat{r}(u,k)}^{u}(x)\ \forall\ k\in K,$ (3)
@@
where $\hat{r}(u,k)$ returns the realization in $\mathcal{R}_{u}=\{1,\ldots,R_{u}\}$ of a random element $u\in\mathcal{U}=\{1,\ldots,U\}$ in scenario $k\in K$. In addition, $\hat{p}_{\hat{r}(u,k)}^{u}(x):\mathcal{X}\rightarrow[0,1]$ is a function on $x$ returning the conditional probability of the realization $\hat{r}(u,k)$ of the random element $u$ in scenario $k\in K$. The function $p_{k}(x)$ can assume other configurations that do not necessarily use $\hat{p}_{\hat{r}(u,k)}^{u}(x)$ but instead return the conditional probability of the whole random vector $\boldsymbol{\xi}$.
A commonly studied case where decisions consist of choosing a probability distribution among a predefined set of alternative distributions is designated the “discrete selection of distributions” (Dupacová 2006, Hellemo et al. 2018) and has been extensively studied. In this case, the decision-dependent probabilities are represented by a probability distribution for each independent random element $u\in\mathcal{U}$ among a set $\mathcal{D}_{u}=\{1,\ldots,D_{u}\}$ of available probability distributions, which could be discrete or continuous. For discrete endogenous random variables, $p_{k}(x)$ are given by
$p_{k}(x)$ $=\prod_{u\in\mathcal{U}}\left(\sum_{d\in\mathcal{D}_{u}}\overline{p}_{\hat{r}(u,k)}^{ud}x_{u}^{d}\right)\ \forall\ k\in K$ (4)
$\sum_{d\in\mathcal{D}_{u}}x_{u}^{d}$ $=1\ \forall\ u\in\mathcal{U},$ (5)
where $x_{u}^{d}\in\{0,1\}$ is a subvector of $x\in\mathcal{X}$ representing the chosen distribution. The parameter $\overline{p}_{\hat{r}(u,k)}^{ud}$ represents the probability of realization $\hat{r}(u,k)$ of random component $u$ considering the selection of distribution $d$, i.e., $\hat{p}_{\hat{r}(u,k)}^{u}(x)=\sum_{d\in D}\overline{p}_{\hat{r}(u,k)}^{ud}x_{u}^{d}$. If a unique distribution in the set $\mathcal{D}=\{1,\ldots,D\}$ is selected for the entire random vector $\boldsymbol{\xi}$ (i.e., for scenario $k\in K$), then $p_{k}(x)=\sum_{d\in\mathcal{D}}\overline{p}^{d}_{k}x^{d}$, where $\overline{p}^{d}_{k}$ is the probability of scenario $k$ according to distribution $d$, and $x^{d}$ are first-stage binary variables equal to one if distribution $d$ is selected and zero otherwise. Finally, the sum of variables $x^{d}$ over $\mathcal{D}$ must also equal one.
Dupacová (2006) also categorizes problems where decisions consist of choosing the parameters of a specific probability distribution as parameter selection. Moreover, Hellemo et al. (2018) define the probability-distortion case, where the chosen probability given $x$ is obtained by combining a set of distributions. We cannot define a general probability mass function $p_{k}(x)$ for these cases because it depends on other assumptions. Nonetheless, we characterize endogenous cases by considering the properties of the probability mass or density function of random vectors, which helps identify appropriate solution methods for each class of problems. We demonstrate this idea in Section 3.

### v4

@@
Solving Programs with Endogenous Uncertainty via Random Variable Transformation
[…]
@@
Real-world decision-making problems often involve decision-dependent uncertainty, where the probability distribution of the random vector depends on the model’s decisions. Few studies focus on two-stage stochastic programs with this type of endogenous uncertainty, and those that do lack general methodologies. We propose a general method for solving a class of these programs based on random variable transformation, a technique widely employed in probability and statistics. The random variable transformation converts a stochastic program with endogenous uncertainty (original program) into an equivalent stochastic program with decision-independent uncertainty (transformed program), for which solution procedures are well-studied. Additionally, endogenous uncertainty usually leads to nonlinear nonconvex programs, which are theoretically intractable. Nonetheless, we show that, for some classical endogenous distributions, the proposed method yields mixed-integer linear or convex programs with exogenous uncertainty. We validate this method by applying it to a network design and facility-protection problem, considering distinct decision-dependent distributions for the random variables. While the original formulation of this problem is nonlinear nonconvex for most endogenous distributions, the proposed method transforms it into mixed-integer linear programs with exogenous uncertainty. We solve these transformed programs with the sample average approximation method. We highlight the superior performance of our approach compared to solving the original program in the case a mixed-integer linear formulation of this program exists.
@@ 1 Introduction
Uncertainty is an inherent characteristic of most real-world decision-making problems (Küçükyavuz and Sen 2017). Stochastic programming is a well-known framework in mathematical programming to model optimization problems with uncertain parameters; see, e.g., (Shapiro et al. 2021) for a summary of the theoretical aspects and the most popular algorithms. Research in this area has focused mainly on exogenous or decision-independent uncertainty (Li and Grossmann 2021), where the random vector is independent of the model decisions. In contrast, endogenous or decision-dependent uncertainty, where the random vector depends on the decisions, has received less attention, although it is prevalent in critical applications such as pre-disaster supply chain management (Peeta et al. 2010, Bhuiyan et al. 2020), and production planning (Ekin 2018).
In the context of endogenous uncertainty, different types of decision dependency may occur (Jonsbrten et al. 1998). Goel and Grossmann (2006) propose classifying decision dependency into two types: For problems with Type 1 endogenous uncertainty or decision-dependent probability, decisions impact the underlying probability distribution of the random parameters. Although little research focuses on this type of uncertainty, as noted by Li and Grossmann (2021), it deals with various problems of practical relevance, such as the design of communication networks, power plants, and warehouses and resource allocation in queuing systems (Pflug 1990). With Type 2 endogenous uncertainty, decisions influence the time at which the realization of the random variables is totally or partially revealed. Thus, with Type 2 endogenous uncertainty, the scenario tree depends on the decisions (Jonsbrten et al. 1998). Problems with Type 2 uncertainty have received more attention than those with Type 1, being mostly multistage programs. Li and Grossmann (2021) provide a concise review of methodologies and algorithms for multistage programs with Type 2 endogenous uncertainty. In the present work, we focus on Type 1 endogenous uncertainty, in particular, on two-stage stochastic mixed-integer linear programs with decision-dependent probability (2-SMIPD), in which the probability distribution depends on the first-stage decisions. In what follows, we refer to Type 1 endogenous uncertainty as endogenous uncertainty.
@@
The main contribution of this work is a general method based on random variable transformation to address 2-SMIPD with discrete or continuous endogenous random variables. This method resolves two main issues in the literature: the lack of generalist methods and the difficulty of solving the usual nonconvex nonlinear programs with endogenous uncertainty. More specifically, the contributions of this work are as follows: (i) We define a transformation function to model the endogenous uncertainty by combining first-stage decisions with decision-independent random variables, thereby converting stochastic programs with endogenous uncertainty (i.e., original programs) into stochastic programs with exogenous uncertainty (i.e., transformed programs). (ii) We identify different techniques to define appropriate transformation functions for endogenous random variables following classical decision-dependent probability distributions, including a commonly studied case known as the “discrete selection of distributions”. (iii) We show that, for these distributions, the obtained transformed programs have a convex continuous relaxation. (iv) We investigate the implications of endogenous uncertainty when applying sampling methods, particularly the sample average approximation (SAA) method, to solve the transformed program. (v) We apply the proposed approach to the network design and facility protection problem (NDFPP), considering different decision-dependent distributions. (vi) Finally, for all endogenous distributions, we analyze the statistical estimators of the SAA applied to the transformed program (SAA-TP) of the NDFPP. In less than one hour on average, the SAA-TP finds feasible solutions with significantly small over-estimated optimality gaps (between $-1\%$ and $1\%$) and a high confidence interval for all decision-dependent distributions. For the NDFPP with most endogenous distributions, the deterministic equivalent of the original program (DE-OP) is nonlinear nonconvex. The exception is the discrete selection of distributions case, for which the DE-OP is a mixed-integer linear program (MILP) obtained by auxiliary linearizations (Bhuiyan et al. 2020). In a considerably shorter time, the SAA-TP obtains a better over-estimator of the optimality gap than the optimality gap obtained by solving this mixed-integer DE-OP (a difference of around $24\%$ on average). Solving the SAA-TP also obtains better solutions than when using an average scenario, i.e., the expected value (EV) problem.
@@
The remainder of this work is structured as follows: Section 2 formalizes the class 2-SMIPD and discusses related works on endogenous uncertainty. Section 3 presents our general methodology to obtain the transformed program and discusses the application of methods such as the SAA to solve this resulting program. Section 4 catalogs techniques to define transformations to various endogenous distributions, and Section 5 explains how to apply this approach to the NDFPP. Finally, Section 5.2 reports the experimental results, comparing the proposed method with other solution procedures, and Section 6 concludes our analyses.
@@ 2 Problem Definition and Related Works
We focus on two-stage stochastic mixed-integer linear programs with decision-dependent probability (2-SMIPD). Let $x\in\mathbb{R}^{n_{1}}_{+}$ be the vector of first-stage variables (i.e., the “here-and-now” decisions) and $\xi(\omega)\in\mathbb{R}^{U}$ a random vector that depends on the random events $\omega\in\Omega$. For the sake of simplicity, we also denote $\bm{\xi}=\xi(\omega)$ and designate $\xi$ as a realization of this random vector. We use $\mathbb{E}_{\bm{\xi}\mid x}[\cdot]$ to denote the expected value operator with respect to $\bm{\xi}$ given the first-stage variables $x$. Finally, $y(\omega)\in\mathbb{R}^{n_{2}}_{+}$ is the vector of second-stage variables given a realization of the random events. The standard mathematical formulation for the 2-SMIPD is:
$\min$ $c^{\top}x+\mathbb{E}_{\bm{\xi}\mid x}\left[Q\bm{(}x,\xi(\omega)\bm{)}\right]$ (1a)
s.t. $Ax\geq b\ ,\ x\in\mathcal{X}$ (1b)
where $\textrm{where }Q\bm{(}x,\xi(\omega)\bm{)}=\min\{q(\omega)^{\top}y(\omega):T(\omega)x+W(\omega)y(\omega)\geq h(\omega),y(\omega)\in\mathcal{Y}\}.$ (1c)
@@
In Program (1), $\mathcal{X}\subseteq\mathbb{R}^{n_{1}}_{+}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n_{2}}_{+}$ are two compact sets imposing non-negativity and continuous, integer, or binary restrictions. The first stage has the matrices $A$, $b$, and $c$ with appropriate dimensions. The second-stage Program (1c) is established after the realization of the random events $\omega\in\Omega$ and has the matrices $q(\omega)$, $h(\omega)$, $T(\omega)$, and $W(\omega)$ with possible random elements and appropriate dimensions. Thus, the random vector $\xi^{T}(\omega)$ combines these random elements, i.e., $\xi^{T}(\omega)=\bm{(}q(\omega)^{\top},h(\omega)^{\top},T(\omega),W(\omega)\bm{)}$, and its support (according to its unconditional distribution on $x$) is denoted by $\Xi\subseteq\mathbb{R}^{U}$. For the sake of clarity, we define the random vectors ${\bm{\xi}_{x}^{\top}=\bm{(}q_{x}(\omega)^{\top},h_{x}(\omega)^{\top},T_{x}(\omega),W_{x}(\omega)\bm{)}}$ for $x\in\mathcal{X}$, which follow the conditional probability distribution of the random vector $\bm{\xi}$ given $x$ and have support $\Xi_{x}\subseteq\Xi$. We also set ${g(x)=\mathbb{E}_{\bm{\xi}\mid x}[Q\bm{(}x,\xi(\omega)\bm{)}]}$.
@@
The 2-SMIPD assumes that the probability distribution of $\bm{\xi}$ depends on the first-stage decisions $x\in\mathcal{X}$ (i.e., it computes the conditional expectation $\mathbb{E}_{\bm{\xi}|x}[\cdot]$ instead of the standard expectation $\mathbb{E}_{\bm{\xi}}[\cdot]$ used in the exogenous case). This endogenous probability influences the properties of the resulting formulation. For example, unlike the exogenous case, the convexity of the continuous relaxation of $Q\bm{(}x,\xi(\omega)\bm{)}$ in Program (1) does not necessarily yield the convexity of the continuous relaxation of $\mathbb{E}_{\bm{\xi}|x}[Q\bm{(}x,\xi(\omega)\bm{)}]$ (Varaiya and Wets 1989, Dupacová 2006). As a result, algorithms commonly used for solving two-stage programs with exogenous uncertainty may not be suitable for 2-SMIPD.
@@
For the sake of simplicity, we next discuss our reasoning assuming that the random vector $\bm{\xi}$ is discrete and has finite support $\Xi$; the process is analogous for the continuous case. We write the DE-OP in its extensive form (i.e., the deterministic equivalent of Program (1)):
@@
where $K$ is the set of realizations of $\bm{\xi}$ (i.e., the set of scenarios defined by the support $\Xi$). For each scenario $k\in K$, $q_{k}$, $h_{k}$, $T_{k}$, and $W_{k}$ have appropriate dimensions, and $\xi_{k}^{\top}=(q_{k}^{\top},h_{k}^{\top},T_{k},W_{k})$. In addition, $p_{k}(x):\mathcal{X}\rightarrow[0,1]$ is a function returning the conditional probability of scenario $k\in K$ given the first-stage variables $x$, in other words, $\bm{\xi}_{x}$ follows the probability distribution defined by $p_{k}(x)$ for every $k\in K$. We remark that scenarios in $\Xi\backslash\Xi_{x}$ have zero probability according to $p_{k}(x)$.
@@ 2.1 Background of Decision-Dependent Probability Distribution
In Program (2), the probability function $p_{k}(x)$ multiplies the second-stage objective term $q_{k}^{\top}y_{k}$, leading to nonlinear terms. Handling this nonlinearity requires describing $p_{k}(x)$, which depends on the assumptions of $\bm{\xi}$. When the $U$ elements of the random vector $\bm{\xi}$ are independent, we have
$p_{k}(x)=\prod_{u=1}^{U}\hat{p}_{\hat{r}(u,k)}^{u}(x)\quad\forall\ k\in K,$ (3)
@@
where $\hat{r}(u,k)$ returns the realization in $\mathcal{R}_{u}=\{1,\ldots,R_{u}\}$ of a random element $u\in\mathcal{U}=\{1,\ldots,U\}$ in scenario $k\in K$. In addition, $\hat{p}_{\hat{r}(u,k)}^{u}(x):\mathcal{X}\rightarrow[0,1]$ is a function on $x$ returning the conditional probability of the realization $\hat{r}(u,k)$ of the random element $u\in\mathcal{U}$ in scenario $k\in K$. The function $p_{k}(x)$ can assume other configurations that do not necessarily use $\hat{p}_{\hat{r}(u,k)}^{u}(x)$ but instead return the conditional probability of the whole random vector $\bm{\xi}$.
A commonly studied case where decisions involve selecting a probability distribution from a predefined set is named “discrete selection of distributions” (Dupacová 2006, Hellemo et al. 2018). In this case, the decision-dependent probability is represented by a probability distribution for each independent random element $u\in\mathcal{U}$ from a set $\mathcal{D}_{u}=\{1,\ldots,D_{u}\}$ of available distributions, which can be either discrete or continuous. For discrete endogenous random variables, we have ${\hat{p}_{\hat{r}(u,k)}^{u}(x)=\sum_{d\in\mathcal{D}_{u}}\overline{p}_{\hat{r}(u,k)}^{ud}x_{u}^{d}}$ for every element $u\in\mathcal{U}$. Here, $x_{u}^{d}\in\{0,1\}$ is a subvector of $x\in\mathcal{X}$ representing the chosen distribution, such that $\sum_{d\in\mathcal{D}_{u}}x_{u}^{d}=1$ for every $u\in\mathcal{U}$. Moreover, the parameter $\overline{p}_{\hat{r}(u,k)}^{ud}$ represents the probability of realization $\hat{r}(u,k)$ of random component $u$ when distribution $d$ is selected. If a distribution from the set $\mathcal{D}=\{1,\ldots,D\}$ is chosen for the entire random vector $\bm{\xi}$ (i.e., for scenario $k\in K$), then $p_{k}(x)=\sum_{d\in\mathcal{D}}\overline{p}^{d}_{k}x^{d}$, where $\overline{p}^{d}_{k}$ is the probability of scenario $k$ under distribution $d$. Similarly, $x^{d}\in\{0,1\}$ are first-stage variables representing the selected distribution, such that $\sum_{d\in\mathcal{D}}x^{d}=1$. Additionally, Dupacová (2006) also categorizes problems involving the selection of parameters for a specific probability distribution as parameter selection. Hellemo et al. (2018) also introduce the probability-distortion case, where the chosen probability given $x$ is derived by combining a set of distributions. A general probability mass function $p_{k}(x)$ cannot be defined for these cases as it relies on additional assumptions.